### v1

# Eigenvalue Tracking of Large-Scale Systems Impacted by Time Delays

Abstract
The paper focuses on tracking eigenvalue trajectories in power system models with time delays. We formulate a continuation-based approach that employs numerical integration to follow eigenvalues as system parameters vary, in the presence of one or multiple delayed variables. The formulation preserves the sparsity of the delay differential-algebraic equation (DDAE) system model and allows treating the delay magnitude itself as a varying parameter with implementation aspects discussed in detail. Accuracy is demonstrated on a modified IEEE 39-bus system with distributed energy resources. Scalability is discussed using a realistic dynamic model of the Irish transmission network.
[…]
This work is supported by Research Ireland (SFI), under grant number 21/SPP/3756 through the NexSys Strategic Partnership Programme.
I Introduction
I-A Motivation
Time delays, arising from the growing volume of data processing and communication in inverter-dominated grids, are an increasingly important factor influencing power system dynamics. In particular, when present in closed-loop control settings, time delays can weaken damping and reduce the system’s stability margin [1, 2, 3, 4, 5]. These effects can be readily assessed using small-signal stability analysis (SSSA), where eigenvalue analysis plays a central role [6]. However, when delays are present, eigenvalue computations and tracking become significantly more complex, particularly in large-scale systems. This paper focuses on continuation-based methods for tracking eigenvalues in power system models with time delays.
I-B Literature Review
Power system dynamics are commonly modeled through a set of differential-algebraic equations [7]. Introducing time delays transforms the model into delay differential-algebraic equations (DDAEs). In the context of SSSA, the stability of linearized DAEs and DDAEs is determined by their eigenvalues, defined as the roots of the system’s characteristic equation. For DDAEs, this equation is transcendental and gives rise to infinitely many eigenvalues [6]. Approximation techniques such as Padé polynomials or spectral discretization yield a finite-dimensional linear eigenvalue problem, which can then be solved using standard numerical algorithms [8]. These include dense-matrix methods (e.g., QR or QZ) for small to medium-sized systems, and sparse techniques (e.g., Krylov) for large-scale cases. However, when tracking eigenvalue trajectories as system parameters vary, repeated use of these solvers can be either computationally expensive or difficult to apply in a robust, automated manner.
Continuation-based methods have been proposed as an alternative framework for tracking eigenvalue trajectories more efficiently. The core idea of continuation methods is to incrementally update the solution of a parameter-dependent problem using previously computed values. In the context of eigenvalue tracking, this solution consists of a subset of the system’s eigenpairs, i.e., eigenvalues with their associated eigenvectors. Setting aside time delays, [9] proposes a continuation-based algorithm that employs Newton’s method to follow eigenvalue trajectories of dynamical systems, highlighting the benefits of including a predictor step before each iteration. The authors in [10] combine this approach with a Cayley transform that updates the dimension of the considered subspace at every iteration, aiming to focus the computation on the modes most critical for stability. To evaluate stability margins of power systems, [11] derives eigenvalue and eigenvector sensitivities with respect to varying parameters and traces their trajectories through numerical integration. This approach makes it possible to identify poorly damped modes, in particular those that are critical for system stability.
Despite the aforementioned advances in continuation-based tracking, only a limited number of works take into account time delays. Among them, [12] uses the derivative of the system’s characteristic equation with respect to system parameters to formulate a matrix equation of Sylvester type. The resulting tangent approximation allows tracking of eigenpairs, but the approach requires elimination of algebraic variables, restricting applicability to small and medium-sized power systems. In contrast, [13] retains the sparsity of the DDAE model. More precisely, [13] uses sensitivities with respect to delays. However, the incremental updates rely solely on Newton iterations, thus being sensitive to discontinuities and poor initialization (e.g., see [14]). Moreover, all existing works exclusively consider constant delays, effectively neglecting real-world communication network effects such as noise and data packet dropouts.
I-C Contribution
This paper develops a continuation-based tracking formulation for power system models described by DDAEs. The formulation is general enough to permit multiple parameters to change simultaneously, as well as treating the delay magnitude as a varying parameter, enabling accurate estimation of stability margins. Extension to account for real-world communication network effects, such as noise and data packet dropouts, is duly discussed. These contributions are demonstrated through a comprehensive case study. First, the method is applied to systems with a single and multiple delays, tracking eigenvalues with respect to parameter variations. Second, the formulation considers the magnitude of the time delay as the varying parameter and its accuracy in capturing the stability margin is demonstrated. Finally, scalability and computational efficiency are assessed on a large-scale dynamic model of the all-island Irish transmission network.
I-D Paper Organization
The remainder of the paper is organized as follows. Section II introduces preliminaries of SSSA. Section III discusses the formulation and implementation aspects of the proposed eigenvalue tracking approach. Section IV presents simulation results, based on the modified 39-bus system and a real-world scale model of the Irish transmission system, highlighting the validity of the method. Finally, Section V draws conclusions and outlines directions for future work.
II Small-Signal Stability Analysis
II-A DDAE Power System Model
Short-term power system stability in the presence of delays can be studied through a set of nonlinear DDAEs [6]:
$\begin{bmatrix}\boldsymbol{T}&\boldsymbol{0}_{n,m}\\ \boldsymbol{R}&\boldsymbol{0}_{m,m}\end{bmatrix}\begin{bmatrix}{\boldsymbol{x}}^{\prime}\\ {\boldsymbol{y}}^{\prime}\end{bmatrix}$ $=\begin{bmatrix}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{x}_{d},\boldsymbol{y}_{d})\\ \boldsymbol{g}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{x}_{d},\boldsymbol{y}_{d})\end{bmatrix}$ (1)
where $\boldsymbol{x}=\boldsymbol{x}(t):[0,\infty)\rightarrow\mathbb{R}^{n}$ are the states of dynamic devices such as generators, dynamic loads, and controllers; $\boldsymbol{y}=\boldsymbol{y}(t):[0,\infty)\rightarrow\mathbb{R}^{m}$ are algebraic variables associated with network equations and auxiliary control setpoints; $\boldsymbol{f}$ and $\boldsymbol{g}$ are nonlinear functions; $\boldsymbol{T}\in\mathbb{R}^{n\times n}$, $\boldsymbol{R}\in\mathbb{R}^{m\times n}$; and $\boldsymbol{0}_{n,m}$ denotes the $n\times m$ zero matrix. The delayed state and algebraic variables are denoted as $\boldsymbol{x}_{d}$ and $\boldsymbol{y}_{d}$, with:
$\boldsymbol{x}_{d}(t)$ $=\{\boldsymbol{x}(t-\tau_{1}),\boldsymbol{x}(t-\tau_{2}),\ldots,\boldsymbol{x}(t-\tau_{\mu})\}$ (2)
$\boldsymbol{y}_{d}(t)$ $=\{\boldsymbol{y}(t-\tau_{1}),\boldsymbol{y}(t-\tau_{2}),\ldots,\boldsymbol{y}(t-\tau_{\mu})\}$
where $\tau_{j}>0$, $j=1,2,\ldots,\mu$, denotes the $j$-th delay and $\mu$ is the total number of delays. For brevity, we denote the $j$-th delayed variable as:
$\boldsymbol{x}^{j}_{d}=\boldsymbol{x}(t-\tau_{j})\ ,\quad\quad j=1,2,\ldots,\mu$ (3)
II-B Eigenvalue Analysis
To study how the system responds to parameter variations, we introduce a scalar continuation parameter $p\in\mathbb{R}$. Then, (1) is rewritten as follows:
$\begin{bmatrix}\boldsymbol{T}(p)&\boldsymbol{0}_{n,m}\\ \boldsymbol{R}(p)&\boldsymbol{0}_{m,m}\end{bmatrix}\begin{bmatrix}{\boldsymbol{x}}^{\prime}\\ {\boldsymbol{y}}^{\prime}\end{bmatrix}$ $=\begin{bmatrix}\boldsymbol{f}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{x}_{d},\boldsymbol{y}_{d},p)\\ \boldsymbol{g}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{x}_{d},\boldsymbol{y}_{d},p)\end{bmatrix}$ (4)
An equilibrium $(\boldsymbol{x}_{o},\boldsymbol{y}_{o}):=[\boldsymbol{x}_{o}^{\intercal},\boldsymbol{y}_{o}^{\intercal}]$ ( indicating the transpose) of (4) is defined assuming the system has been at rest for time equal or larger than the maximum delay. Considering small disturbances, (4) can be linearized around the equilibrium, as follows:
$\begin{bmatrix}\boldsymbol{T}(p)&\boldsymbol{0}_{n,m}\\ \boldsymbol{R}(p)&\boldsymbol{0}_{m,m}\end{bmatrix}\begin{bmatrix}\Delta{\boldsymbol{x}^{\prime}}\\ \Delta{\boldsymbol{y}^{\prime}}\end{bmatrix}=\begin{bmatrix}\boldsymbol{{f}}_{\hskip-0.56905ptx}(p)&\boldsymbol{{f}}_{\hskip-0.56905pty}(p)\\ \boldsymbol{{g}}_{\hskip-0.56905ptx}(p)&\boldsymbol{{g}}_{\hskip-0.56905pty}(p)\end{bmatrix}\begin{bmatrix}\Delta{\boldsymbol{x}}\\ \Delta{\boldsymbol{y}}\end{bmatrix}$
[…]
where $\Delta\boldsymbol{x}=\boldsymbol{x}-\boldsymbol{x}_{o}$, $\Delta\boldsymbol{y}=\boldsymbol{y}-\boldsymbol{y}_{o}$; $\boldsymbol{{f}}_{\hskip-0.56905ptx}$, $\boldsymbol{{f}}_{\hskip-0.56905pty}$, $\boldsymbol{{g}}_{\hskip-0.56905ptx}$, $\boldsymbol{{g}}_{\hskip-0.56905pty}$ and $\boldsymbol{{f}}_{\hskip-0.56905ptx,j}$, $\boldsymbol{{f}}_{\hskip-0.56905pty,j}$, $\boldsymbol{{g}}_{\hskip-0.56905ptx,j}$, $\boldsymbol{{g}}_{\hskip-0.56905pty,j}$ are the delay-free and delayed Jacobians, respectively, evaluated at $(\boldsymbol{x}_{o},\boldsymbol{y}_{o})$. Equation (5) is of the form:
$\boldsymbol{\rm E}(p)\,\mbox{$\boldsymbol{\rm x}$}^{\prime}=\boldsymbol{\rm A}_{0}(p)\,\mbox{$\boldsymbol{\rm x}$}+\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}(p)\,\mbox{$\boldsymbol{\rm x}$}^{j}_{d}$ (6)
where $\mbox{$\boldsymbol{\rm x}$}=(\Delta\boldsymbol{x},\Delta\boldsymbol{y})$, $\mbox{$\boldsymbol{\rm x}$}^{j}_{d}=(\Delta\boldsymbol{x}^{j}_{d},\Delta\boldsymbol{y}^{j}_{d})$ and
$\boldsymbol{\rm E}(p)$ $\equiv\begin{bmatrix}\boldsymbol{T}(p)&\boldsymbol{0}_{n,m}\\ \boldsymbol{R}(p)&\boldsymbol{0}_{m,m}\\ \end{bmatrix}\,,\ \boldsymbol{\rm A}_{0}(p)\equiv\begin{bmatrix}\boldsymbol{{f}}_{\hskip-0.56905ptx}(p)&\boldsymbol{{f}}_{\hskip-0.56905pty}(p)\\ \boldsymbol{{g}}_{\hskip-0.56905ptx}(p)&\boldsymbol{{g}}_{\hskip-0.56905pty}(p)\\ \end{bmatrix}$ (7)
$\boldsymbol{\rm A}_{j}(p)$ $\equiv\begin{bmatrix}\boldsymbol{{f}}_{\hskip-0.56905ptx,j}(p)&\boldsymbol{{f}}_{\hskip-0.56905pty,j}(p)\\ \boldsymbol{{g}}_{\hskip-0.56905ptx,j}(p)&\boldsymbol{{g}}_{\hskip-0.56905pty,j}(p)\\ \end{bmatrix}\,$
Applying the Laplace transform to (6):
$(s\boldsymbol{\rm E}(p)-\boldsymbol{\rm A}_{0}(p)-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}(p)e^{-s\tau_{j}})\;\mathcal{L}\{\mbox{$\boldsymbol{\rm x}$}\}=\boldsymbol{\rm E}(p)\mbox{$\boldsymbol{\rm x}$}(0)$ (8)
where $s$ is a complex variable in the $S$-plane. The associated eigenvalue problem is:
${\boldsymbol{\rm P}}(s,p)\;\boldsymbol{\rm\phi}$ $=\boldsymbol{0}_{r,1}\;,\quad r=n+m$ (9)
where any value of $s$ that satisfies (9) is an eigenvalue of the matrix function:
${\boldsymbol{\rm P}}(s,p)=s\boldsymbol{\rm E}(p)-\boldsymbol{\rm A}_{0}(p)-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}(p)e^{-s\tau_{j}}$ (10)
with $\boldsymbol{\phi}$ being the associated right eigenvector.
III Eigenvalue tracking
This section presents a continuation-based approach for tracking the eigenvalues of power systems with time delays. The method is first introduced for the case of a single delayed variable, then generalized to systems with multiple delays, and finally extended to allow the delay magnitude itself to vary as the continuation parameter.
III-A Tracking in the Presence of a Single Delay
We begin with the case of a single delay $\tau=\tau_{1}$ ($\mu=1$). In this case, (10) takes the form:
${\boldsymbol{\rm P}}=s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{1}e^{-s\tau}$ (11)
where the dependence on $p$ is omitted for brevity. Differentiation of (9) with respect to $p$ gives:
$\dot{\boldsymbol{\rm P}}\;\boldsymbol{\rm\phi}+{\boldsymbol{\rm P}}\;\dot{\boldsymbol{\rm\phi}}$ $=\boldsymbol{0}_{r,1}$ (12)
where
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm A}}_{1}e^{-{s}\tau}+\tau\dot{s}\boldsymbol{\rm A}_{1}e^{-{s}\tau}$ (13)
with $\dot{s}=\partial s/\partial p$, $\dot{\boldsymbol{\rm E}}=\partial{\boldsymbol{\rm E}}/\partial p$, $\dot{\boldsymbol{\rm A}}_{0}=\partial{\boldsymbol{\rm A}}_{0}/\partial p$, $\dot{\boldsymbol{\rm A}}_{1}=\partial{\boldsymbol{\rm A}}_{1}/\partial p$, $\dot{\boldsymbol{\rm\phi}}=\partial{\boldsymbol{\rm\phi}}/\partial p$. Equivalently, equation (12) is:
$(\boldsymbol{\rm E}+\tau\boldsymbol{\rm A}_{1}e^{-s\tau})\boldsymbol{\rm\phi}\dot{s}+(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{1}e^{-s\tau})\dot{\boldsymbol{\rm\phi}}=$ (14)
$-(s\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm A}}_{1}e^{-s\tau})\boldsymbol{\rm\phi}$
System (14), which describes the evolution of a single eigenpair with respect to $p$, consists of $r$ equations and $r+1$ unknowns, $s$ and $\boldsymbol{\phi}$. To make this system well-posed, it is closed by imposing the following eigenvector normalization:
$\boldsymbol{\phi}^{\intercal}\boldsymbol{\phi}=c$ (15)
where $c$ is an imposed constant, e.g., $c=1$. Differentiation of (15) gives:
$\boldsymbol{\phi}^{\intercal}\dot{\boldsymbol{\phi}}$ $=0$ (16)
Combining (14) and (16):
$\begin{bmatrix}s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{1}e^{-s\tau}&(\boldsymbol{\rm E}+\tau\boldsymbol{\rm A}_{1}e^{-s\tau})\boldsymbol{\rm\phi}\\ {\boldsymbol{\rm\phi}}^{\intercal}&0\end{bmatrix}\begin{bmatrix}\dot{\boldsymbol{\rm\phi}}\\ \dot{s}\end{bmatrix}$=
$\begin{bmatrix}-({s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm A}}_{1}e^{-s\tau}){\boldsymbol{\rm\phi}}\\ 0\end{bmatrix}$ (17)
Splitting real and imaginary parts, i.e., $\boldsymbol{\rm\phi}=\boldsymbol{\rm\phi}_{\rm r}+\jmath\boldsymbol{\rm\phi}_{\rm i}$, $s=s_{\rm r}+\jmath s_{\rm i}$, and by setting $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}$, $\boldsymbol{\rm A}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}$, $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\cos{(s_{\rm i}\tau)}=\boldsymbol{\rm C}_{\scriptscriptstyle D}$ and $\dot{\boldsymbol{\rm A}}_{1}e^{-s_{\rm r}\tau}\sin{(s_{\rm i}\tau)}=\boldsymbol{\rm S}_{\scriptscriptstyle D}$ we arrive at the following system:
$\boldsymbol{\rm M}(\boldsymbol{\rm y})\;\dot{\boldsymbol{\rm y}}=\boldsymbol{\rm h}(\boldsymbol{\rm y})$ (18)
where $\boldsymbol{\rm y}=\boldsymbol{\rm y}(p)=(\boldsymbol{\rm\phi}_{\rm r},\boldsymbol{\rm\phi}_{\rm i},s_{\rm r},s_{\rm i})$, with $\boldsymbol{\rm y}\in\mathbb{R}^{2r+2}$; with:
$\boldsymbol{\rm M}(\boldsymbol{\rm y})$ $=\begin{bmatrix}\boldsymbol{\rm M}_{1}&\boldsymbol{\rm M}_{2}\\ \boldsymbol{\rm M}_{3}&\boldsymbol{\rm 0}_{2,2}\end{bmatrix}$
The quantities $\boldsymbol{\rm h}$, $\boldsymbol{\rm M}_{1}$, $\boldsymbol{\rm M}_{2}$, $\boldsymbol{\rm M}_{3}$ are in this case defined as:
$\boldsymbol{\rm h}$ $=\begin{bmatrix}(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm C}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm r}+(s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm S}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm i}\\ -(s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm S}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm r}+(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm C}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$
$\boldsymbol{\rm M}_{1}$ $=\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm C}&-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm S}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm S}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm C}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\begin{bmatrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})&\!\!-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!-\!\tau(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})\\ \boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!+\!\tau(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})&\!\!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})\end{bmatrix}$
$\boldsymbol{\rm M}_{3}$ $=\begin{bmatrix}\boldsymbol{\phi}_{\rm r}^{\intercal}&-\boldsymbol{\rm\phi}_{\rm i}^{\intercal}\\ \boldsymbol{\rm\phi}_{\rm i}^{\intercal}&\boldsymbol{\phi}_{\rm r}^{\intercal}\end{bmatrix}\hskip 139.4185pt$
III-B Generalization to Multiple Delays
When multiple delays are present in the system, the matrix function $\boldsymbol{\rm P}$ is given by (10). Differentiating it with respect to the continuation parameter $p$ gives:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\sum_{j=1}^{\mu}(\dot{\boldsymbol{\rm A}}_{j}e^{-{s}\tau_{j}}-\tau_{j}\dot{s}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})$ (19)
Substituting into (12) yields:
$\hskip-5.69054pt(\boldsymbol{\rm E}+\sum_{j=1}^{\mu}\tau_{j}$ $\boldsymbol{\rm A}_{j}e^{-s\tau_{j}})\boldsymbol{\rm\phi}\dot{s}+({s}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})\dot{\boldsymbol{\rm\phi}}$
$=-({s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\sum_{j=1}^{\mu}\dot{\boldsymbol{\rm A}}_{j}e^{-s\tau_{j}})\boldsymbol{\rm\phi}$ (20)
Combining with (16) and splitting eigenvalue and eigenvector into real and imaginary parts leads to a system in the form of (18), where $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ are defined to account for all $\mu$ delays. Their detailed definitions are provided in the Appendix.
III-C Delay Magnitude as Varying Parameter
In this subsection, we track the evolution of eigenvalues as the delay magnitude varies. We first consider the case where the system contains a single delayed variable, whose time delay has a varying magnitude and will be denoted by $\tau_{\ell}$. Its corresponding delayed state matrix is denoted by $\boldsymbol{\rm A}_{\ell}$. Given $p=\tau_{\ell}$, the matrix function $\boldsymbol{\rm P}$} is:
${\boldsymbol{\rm P}}=s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-{s}p}$ (21)
Differentiation of (21) with respect to $p$ gives:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm A}}_{\ell}e^{-{s}p}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$
In the above expression, $\boldsymbol{\rm A}_{\ell}$ does not explicitly depend on $\tau$, so $\dot{\boldsymbol{\rm A}}_{\ell}=0$. Substituting into (12) yields:
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-sp})\dot{\boldsymbol{\rm\phi}}+(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}pe^{-sp})\boldsymbol{\rm\phi}\dot{s}$ $=$
$-s\dot{\boldsymbol{\rm E}}\,\boldsymbol{\rm\phi}+\dot{\boldsymbol{\rm A}}_{0}\boldsymbol{\rm\phi}-\boldsymbol{\rm A}_{\ell}e^{-sp}s$ $\boldsymbol{\rm\phi}$ (22)
Splitting real and imaginary parts of (22), we arrive at a system in the form of (18), where $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ are in this case functions of the varying delay magnitude $p=\tau_{\ell}$. Their definitions are detailed in the Appendix.
We extend the above formulation to systems with multiple delays, treating one of them as the continuation parameter, i.e., $\tau_{\ell}=p$. The definition of $\boldsymbol{\rm P}$ in this case is as follows:
$\boldsymbol{\rm P}=s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}}-\boldsymbol{\rm A}_{\ell}e^{-{s}p}$ (23)
We note that the term corresponding to $\tau_{\ell}$ remains separate, even if $\tau_{\ell}$ equals some $\tau_{j},\;j=1,2,\ldots,\mu$. Given that the delayed state matrices $\boldsymbol{\rm A}_{\ell}$ and $\boldsymbol{\rm A}_{j}$ do not explicitly depend on $\tau_{\ell}$, the derivative of (23) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}+\sum_{j=1}^{\mu}\tau_{j}\dot{s}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}}+\boldsymbol{\rm A}_{\ell}(\dot{s}p+s)e^{-{s}p}$ (24)
and (12) is equivalently written as:
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}e^{-{s}p}-\sum_{j=1}^{\mu}\boldsymbol{\rm A}_{j}e^{-{s}\tau_{j}})\dot{\boldsymbol{\rm\phi}}$
$+(\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}\,p\,e^{-sp}+\sum_{j=1}^{\mu}\tau_{j}\boldsymbol{\rm A}_{j}e^{-s\tau_{j}})\boldsymbol{\rm\phi}\dot{s}=$
$-({s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm A}_{\ell}e^{-sp}s)\boldsymbol{\rm\phi}$ (25)
By splitting eigenvalue and eigenvector into real and imaginary parts we arrive at a system in the form of (18). The derivation of $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ is provided in the Appendix.
III-D Communication Delays with Noise and Data Dropouts
The derivations above consider constant delays. In this section, we discuss how they can be conveniently generalized to capture the effects of realistic wide area measurement system (WAMS) latencies with noise and data packet dropouts. To this end, we consider the composite time-varying delay model proposed in [1], wherein for the purpose of SSSA, packet dropouts and noise are captured through the functions $h_{p}(s)$ and $h_{s}(s)$, respectively:
$h_{p}(s)=\frac{1-p_{dr}}{s}[1+(p_{dr}-1)\frac{e^{-sT}}{1-p_{dr}e^{-sT}}]$ (26)
$h_{s}(s)=(1+\frac{\alpha}{1-p_{dr}}s)^{-b}$ (27)
where $\alpha$ and $b$ are the scale and shape factor of the Gamma distribution; $p_{dr}$ is the packet dropout rate; and $T$ is the normal delivery period for each data packet.
Equation (10) takes the adjusted form:
${\boldsymbol{\rm P}}=s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-h_{p}(s)h_{s}(s)\boldsymbol{\rm A}_{1}e^{-s\tau_{0}}$ (28)
where $\tau_{0}$ is the constant component of the WAMS delay. By setting $\boldsymbol{\rm S}_{T}=h_{p}(s)h_{s}(s)\boldsymbol{\rm A}_{1}e^{-s\tau_{0}}$, the derivative of (28) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm S}}_{T}$ (29)
where
$\dot{\boldsymbol{\rm S}}_{T}=$ $(\dot{\boldsymbol{\rm A}}_{1}h_{p}h_{s}+\boldsymbol{\rm A}_{1}(\frac{\partial{h_{p}}}{\partial{s}}h_{s}+h_{p}\frac{\partial{h_{s}}}{\partial{s}}))e^{-s\tau_{0}}$ (30)
$-\tau_{0}h_{p}h_{s}\boldsymbol{\rm A}_{1}e^{-s\tau_{0}}\dot{s}$
$=$ $\boldsymbol{\rm S}_{TD}-\tau_{0}\boldsymbol{\rm S}_{T}\dot{s}$
We note that in the case of a constant delay, i.e., $\tau(t)=\tau_{0}$, it is trivial to deduce from [1] that $h_{p}(s)=h_{s}(s)=1$ and $\partial{h_{p}}/\partial{s}=\partial{h_{s}}/\partial{s}=0$; therefore, equations (28) and (29) are simplified to (11) and (13) respectively.
Substituting (28) and (29) into (12) yields:
$(\boldsymbol{\rm E}+\tau_{0}\boldsymbol{\rm S}_{T})\boldsymbol{\rm\phi}\dot{s}+(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T})\dot{\boldsymbol{\rm\phi}}$ $=$ (31)
$-(s\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\boldsymbol{\rm S}_{TD})\boldsymbol{\rm\phi}$
By splitting real and imaginary parts of (31), we arrive at a system in the form of (18), where the definitions of $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ are detailed in the Appendix.
III-E Numerical Integration
Given a parameter range $p_{init},p_{fin}]$, tracking is performed by numerically integrating (18). For the initial value $p_{init}$, the state vector $\boldsymbol{\rm y}(p_{init})$ is obtained by solving the eigenvalue problem (9). An effective method, in our experience, is to transform the linearized DDAEs into an equivalent system of partial differential equations, and then reduce it through Chebyshev polynomials to a finite-dimensional linear eigenvalue problem. The spectral discretization technique used in this paper is described in [15]. In subsequent steps, if the examined variation alters the power flow solution, the latter is recomputed. Otherwise, if the varying parameters correspond to dynamic devices, the last step is skipped; the operating point is updated from the steady-state DDAEs and the matrices $\boldsymbol{\rm A}_{0}$, $\boldsymbol{\rm A}_{1}$ and $\boldsymbol{\rm E}$ are reconstructed. Matrix derivatives are computed numerically using first-order finite differences.
Integration proceeds iteratively until $p=p_{fin}$; firstly, with step size $\Delta p$, the parameter is updated as $p_{k+1}=p_{k}+\Delta p$. Then, the mass matrix $\boldsymbol{\rm M}(\boldsymbol{\rm y}_{k})$ and right-hand side $\boldsymbol{\rm h}(\boldsymbol{\rm y}_{k})$ are evaluated, and the state vector is advanced using a numerical integration method. For instance, using forward Euler gives:
$\boldsymbol{\rm y}_{k+1}=\boldsymbol{\rm y}_{k}+\Delta p\,\boldsymbol{\rm M}^{-1}(\boldsymbol{\rm y}_{k})\,\boldsymbol{\rm h}(\boldsymbol{\rm y}_{k})$ (32)
The product $\boldsymbol{\rm M}^{-1}(\boldsymbol{\rm y}_{k})\,\boldsymbol{\rm h}(\boldsymbol{\rm y}_{k})$ can be determined through LU decomposition of $\boldsymbol{\rm M}(\boldsymbol{\rm y}_{k})$.
We note that numerical integration schemes inherently introduce discretization error, which depends on the step size. This error can be practically eliminated by adding a Newton-based corrector step that recomputes the solution to within a prescribed tolerance, at additional computational cost.
IV Case Study
In this section, we apply the continuation-based eigenvalue tracking formulation of Section III to two power system models. First, we assess accuracy based on a modified version of the IEEE 39-bus system [16]. We then discuss scalability based on a large-scale model of the Irish transmission system.
IV-A 39-Bus System
We consider a modified version of the IEEE 39-bus system, wherein the machines at buses 30, 34, 35, 36, 37 are replaced by distributed energy resources of equal capacity. Each DER is modeled with an inner current control loop and outer voltage and frequency regulation loops [17].
(a) FR mode: increasing the DER frequency controllers’ droop constant.
(b) Critical electromechanical mode: Reducing the inertia of the SM at bus 39.
Figure 1: 39-bus system: tracking in the presence of a single delay.
We begin by demonstrating the proposed tracking method on the system’s frequency regulation (FR) mode, also referred to as common low-frequency mode in the literature [18, 19]. This is a system-wide mode originating from the dynamics of frequency control loops [14]. For the examined system, the natural frequency of the FR is initially $0.005$ Hz. To illustrate the process, the frequency control droop constant $\rm R_{DER}$ of the five DERs is varied from $0.05$ to $0.5$. A single delay of $\tau=0.01$ s is introduced in the frequency signal of the power system stabilizer (PSS) connected to the synchronous machine (SM) at bus 31. The tracked eigenvalue, initially at $0.0276+\jmath 0.0165$, follows the trajectory shown in Fig. 1(a). Tracking based on (18) accurately captures the trajectory and shows how the oscillatory behavior of the FR mode is progressively suppressed as the relative share of frequency control shifts from the DERs to the SMs. At the point where the complex pair coalesces on the real axis, the eigenvalue becomes defective with algebraic multiplicity $2$ and geometric multiplicity $1$. This corresponds to a simple quadratic fold [20, 14]. As $\rm R_{DER}$ is further increased, the defective eigenvalue splits into two distinct real eigenvalues. The tracking method inherently follows one of the two emerging branches, as determined by the evolution of (18). In this example, the eigenvalue on the left branch of the fold is traced, and Fig. 1(a) shows two additional folds occurring along this trajectory. In practice, when such splitting occurs, the process is reinitialized by recomputing the eigendecomposition of $\boldsymbol{\rm P}$ and selecting the eigenpair corresponding to the branch of interest.
As a second example, the inertia constant of the system’s largest SM is gradually reduced from $50$ to $5$. Figure 1(b) illustrates the resulting trajectory of the system’s least damped mode, initially located at $0.4745\pm\jmath 8.8572$ and corresponding to the local electromechanical mode of the SM at bus 31. Tracking based on (18) shows how the eigenvalue moves as inertia decreases, eventually crossing into the unstable region.
(a) FR mode.
(b) Inter-machine oscillation mode.
Figure 2: 39-bus system: tracking in the presence of multiple delays.
We next assess the method in the presence of multiple delays. To this end, we introduce delays to the input signals of the system’s automatic voltage regulators, PSSs and DER frequency and voltage controllers, yielding a total of 20 time-delayed variables. The resulting $\boldsymbol{\rm P}$ has dimension $558\times 558$. For each delay $\tau_{j}$, the spectral discretization in [15] uses $N$ Chebyshev nodes in the interval $\tau_{j},0]$, which increases the size of $\boldsymbol{\rm P}$ by $N+1$. Thus, considering $N=10$, the dimension of the approximated linear eigenvalue problem grows by $220$ compared to the delay-free model. We then trace how variations in the SM droop constants affect the eigenvalues corresponding to (i) the system’s FR mode and (ii) the inter-machine mode with the largest participation in the rotor speeds of the SMs at buses $32$ and $39$, initially located at $0.4425\pm\jmath 0.2052$. The calculated trajectories, shown in Fig. 2 confirm that the proposed tracking method accurately captures eigenvalue trajectories in the presence of multiple delays.
We next examine the tracking accuracy when the delay magnitude is treated as the continuation parameter, as described in Section III-C. The analysis focuses on the local electromechanical mode of the SM at bus 31. The delay $\tau_{\rm PSS}$ in the input signal of the corresponding PSS is varied from $0.01$ to $0.6$ s, producing the trajectory shown in Fig. 3(a). As $\tau_{\rm PSS}$ increases, the traced eigenvalue becomes progressively less damped and eventually crosses into instability for $\tau_{\rm PSS}>0.1$ s. Further increasing $\tau_{\rm PSS}$ beyond $0.28$ s reverses this trend and eventually leads the system to regain small-signal stability for $\tau_{\rm PSS}>0.5$ s. This behavior is further validated via a time-domain simulation, considering a three-phase fault at bus 6, cleared after 80 ms by opening line 5–6. The response of the center-of-inertia (COI) frequency is shown in Fig. 3(b) and confirms the conclusions drawn by the eigenvalue tracking.
(a) Tracing local electromechanical mode as delay is increased.
(b) Time-domain simulation for different delay values.
Figure 3: 39-bus system, SM at bus 31: PSS with input signal delay.
IV-B Irish System
We next assess the scalability of the proposed eigenvalue tracking method using a 1,502-bus dynamic model of the All-Island Irish Transmission System (AIITS). The delay-free DAE model has 1,629 states and 9,897 algebraic variables. Its dimension, $11,526\times 11,526$, renders the use of standard, dense QZ-based eigensolvers impractical. We first examine the accuracy of (18) in tracing poorly damped electromechanical modes of the AIITS. The DDAE model includes 28 delayed variables, affecting voltage and frequency measurements of the system’s AVRs and PSSs. The delays, ranging from $0.009$ to $0.022$ s, increase the dimension of the approximated linear eigenvalue problem to $11,834\times 11,834$. The effect of raising the gains of the system’s PSSs from $1.5$ to $10$, is shown in Fig. 4. Although the examined eigenvalues are tightly clustered in the complex plane, the proposed approach efficiently traces all of them.
Figure 4: AIITS: tracking poorly damped electromechanical modes in the presence of multiple delays.
We next apply the formulation in Section III-C to trace the trajectories of the system’s poorly damped electromechanical modes, when the time delay of the frequency signal of the PSS, at bus 717 increases from $0.01$ to $0.5$ s. The results, displayed in Fig. 5, confirm the accuracy of the proposed approach.
Figure 5: AIITS, SM at bus 717: increasing PSS input signal delay.
V Conclusion
This paper proposes a continuation-based eigenvalue tracking technique for power systems impacted by time delays. The formulation retains the sparsity of the system’s DDAE model and considers a continuation parameter which enables multiple system properties – including delay magnitudes – to be expressed as functions of it and varied simultaneously. Case studies demonstrate the accuracy and computational efficiency of the proposed approach. Future work will focus on applying the proposed approach to guide the design of delay-robust control schemes for DERs.
In this section, we derive the expressions of $\boldsymbol{\rm M}$ and $\boldsymbol{\rm h}$ for the cases considered in Sections III-B, III-CIII-D, i.e., for systems with multiple delays, systems including a varying delay and systems including a WAMS delay.
-A Systems with Multiple Delays
We define the functions $h_{r}(t)=e^{-s_{\rm r}t}\cos{(s_{\rm i}t)}$ and $h_{i}(t)=e^{-s_{\rm r}t}\sin{(s_{\rm i}t)}$ and consider the notation $h^{j}_{r}=h_{r}(\tau_{j})$ and $h^{j}_{i}=h_{i}(\tau_{j})$. By splitting real and imaginary parts of (20) and setting $\boldsymbol{\rm C}=[\boldsymbol{\rm A}_{j}h^{j}_{r}]^{\intercal}$, $\boldsymbol{\rm S}=[\boldsymbol{\rm A}_{j}h^{j}_{i}]^{\intercal}$, $\boldsymbol{\rm C}_{\scriptscriptstyle D}=[\dot{\boldsymbol{\rm A}}_{j}h^{j}_{r}]^{\intercal}$, $\boldsymbol{\rm S}_{\scriptscriptstyle D}=[\dot{\boldsymbol{\rm A}}_{j}h^{j}_{i}]^{\intercal}$, we obtain a system of the form (18), where
$\boldsymbol{\rm h}$ $=\begin{bmatrix}(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}+\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm r}+(s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm i}\\ (-s_{\rm i}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm r}+(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}+\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}_{\scriptscriptstyle D})\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$
$\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm M}_{1}$ $=\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}&-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\begin{bmatrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})&\!\!-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!-\!\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})\!\\ \boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!+\!\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})&\!\!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})\end{bmatrix}$
with $\boldsymbol{\rm J}_{\mu}=[\boldsymbol{I}_{r}\;\boldsymbol{I}_{r}\;\ldots\;\boldsymbol{I}_{r}]\in\mathbb{R}^{r\times\mu r}$ and $\boldsymbol{\rm J}_{\tau\mu}=[\tau_{1}\boldsymbol{I}_{r}\;\tau_{2}\boldsymbol{I}_{r}\;\ldots\;\tau_{\mu}\boldsymbol{I}_{r}]\in\mathbb{R}^{r\times\mu r}$; $\boldsymbol{I}_{r}$ is the $r\times r$ identity matrix.
-B Systems with Delay as Varying Parameter
Splitting real and imaginary parts of (22) and by considering the functions $h_{r}$ and $h_{i}$ defined above for $\tau_{\ell}=p$, we arrive at the form of (18), where in this case:
$\boldsymbol{\rm h}=\begin{bmatrix}\boldsymbol{\rm h}_{1}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm h}_{2}\boldsymbol{\phi}_{\rm i}\\ -\boldsymbol{\rm h}_{2}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm h}_{1}\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$ (33)
with
$\;\;\,\boldsymbol{\rm h}_{1}=$ $\;\dot{\boldsymbol{\rm A}}_{0}-s_{\rm r}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\;\;\,\boldsymbol{\rm h}_{2}=$ $\;s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
$\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm M}_{1}=\!$ $\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}&-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}\!=\!\!$ $\begin{bmatrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!p\boldsymbol{\rm A}_{\ell}(\boldsymbol{\phi}_{\rm r}h^{\ell}_{r}\!+\!\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{i})&\!\!-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!+\!p\boldsymbol{\rm A}_{\ell}(\boldsymbol{\phi}_{\rm r}h^{\ell}_{i}\!-\!\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{r})\!\\ \!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!-\!p\boldsymbol{\rm A}_{\ell}(\boldsymbol{\phi}_{\rm r}h^{\ell}_{i}\!-\!\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{r})\!&\!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!p\boldsymbol{\rm A}_{\ell}(\boldsymbol{\phi}_{\rm r}h^{\ell}_{r}\!+\!\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{i})\!\end{bmatrix}$
If the system includes multiple delays, then we also consider the matrices $\boldsymbol{\rm C}$, $\boldsymbol{\rm S}$, $\boldsymbol{\rm J}_{\mu}$ and $\boldsymbol{\rm J}_{\tau\mu}$. Splitting real and imaginary parts of (25) we arrive at the form of (18), with $\boldsymbol{\rm h}$ given by (33) and where in this case
$\boldsymbol{\rm h}_{1}=$ $\dot{\boldsymbol{\rm A}}_{0}-s_{\rm r}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\boldsymbol{\rm h}_{2}=$ $s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
$\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm M}_{1}$ $=\!\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}\!-\!\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}\!-\!\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}\!\!&\!\!-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}+\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm S}\!\!&\!\!s_{\rm r}\boldsymbol{\rm E}\!-\!\boldsymbol{\rm A}_{0}\!-\!\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}-\boldsymbol{\rm J}_{\mu}\boldsymbol{\rm C}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\left[\begin{matrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})+\boldsymbol{\rm A}_{\ell}p(\boldsymbol{\phi}_{\rm r}h^{\ell}_{r}+\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{i})&\!\!\!\!\ldots\\ \boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}+\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm i}-\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm r})-\boldsymbol{\rm A}_{\ell}p(\boldsymbol{\phi}_{\rm r}h^{\ell}_{i}-\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{r})$$\!\!\ldots\end{matrix}\right$.
$\left.\begin{matrix}\ldots\!\!\!&-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}-\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm i}-\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm r})+\boldsymbol{\rm A}_{\ell}p(\boldsymbol{\phi}_{\rm r}h^{\ell}_{i}-\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{r})\\ \ldots\!\!\!&\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm J}_{\tau\mu}(\boldsymbol{\rm C}\boldsymbol{\phi}_{\rm r}+\boldsymbol{\rm S}\boldsymbol{\phi}_{\rm i})+\boldsymbol{\rm A}_{\ell}p(\boldsymbol{\phi}_{\rm r}h^{\ell}_{r}+\boldsymbol{\rm\phi}_{\rm i}h^{\ell}_{i})\end{matrix}\right]$
-C Communication Delays with Noise and Data Dropouts
We set $\boldsymbol{\rm S}_{T}^{r}=\Re\{\boldsymbol{\rm S}_{T}\}$, $\boldsymbol{\rm S}_{T}^{i}=\Im\{\boldsymbol{\rm S}_{T}\}$, $\boldsymbol{\rm S}_{TD}^{r}=\Re\{\boldsymbol{\rm S}_{TD}\}$ and $\boldsymbol{\rm S}_{TD}^{i}=\Im\{\boldsymbol{\rm S}_{TD}\}$. Splitting real and imaginary parts of (31) leads to a system in the form of (18), where in this case $\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm h}$ $=\begin{bmatrix}(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm S}_{TD}^{r})\boldsymbol{\phi}_{\rm r}+(s_{\rm i}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm S}_{TD}^{i})\boldsymbol{\phi}_{\rm i}\\ -(s_{\rm i}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm S}_{TD}^{i})\boldsymbol{\phi}_{\rm r}+(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm S}_{TD}^{r})\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$
$\boldsymbol{\rm M}_{1}$ $=\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T}^{r}&-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm S}_{T}^{i}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm S}_{T}^{i}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T}^{r}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\begin{bmatrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm i})&\!\!-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!-\!\tau_{0}(\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm i})\\ \boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm i})&\!\!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm i})\end{bmatrix}$